please show all the steps ?
all the way to the solution please

QUESTION 5 Consider the IVP: x + (x2 +4tx=0, x(0) = 1, x(0) = 2 the nearest thousandth (3 decimal places). In the answer U
Answer #1


Given Ivp is x+ (0)2+4+x=0; x 10) = | - x(0)=2 step length At = h = 0.5 o 0.5 lesbian put nay y = =-(2) 4th E-y*- 4n & x= & Y = –2-105-12 Forf in 7 (1)= = -1 & 2 (1) = 8 = 12 Similarly sen i = 2 = Mg + f (mg, Yz, to) = ng tya 2.4+1-12) = -13 73 -